The following table contains the demand from the last 10 months: MONTH ACTUAL DEMAND 1 33 2 32 3 31 4 38 5 42 6 38 7 42 8 42 9 43 10 43 a. Calculate the single exponential smoothing forecast for these data using an α of 0.20 and an initial forecast (F1) of 33. (Round your intermediate calculations and answers to 2 decimal places.) Month Exponential Smoothing 1 2 3 4 5 6 7 8 9 10
Answer #1

F(t) = F(t-1) + (Alpha * (A(t-1) - F(t-1)))

Where F(t-1) is the forecast for the previous period and A(t-1) is the actual demand for the previous period.

FORECAST 2 = 33 + (0.2 * (33 - 33)) = 33

FORECAST 3 = 33 + (0.2 * (32 - 33)) = 32.8

FORECAST 4 = 32.8 + (0.2 * (31 - 32.8)) = 32.44

FORECAST 5 = 32.44 + (0.2 * (38 - 32.44)) = 33.55

FORECAST 6 = 33.55 + (0.2 * (42 - 33.55)) = 35.24

FORECAST 7 = 35.24 + (0.2 * (38 - 35.24)) = 35.79

FORECAST 8 = 35.79 + (0.2 * (42 - 35.79)) = 37.03

FORECAST 9 = 37.03 + (0.2 * (42 - 37.03)) = 38.02

FORECAST 10 = 38.02 + (0.2 * (43 - 38.02)) = 39.02

FORECAST 11 = 39.02 + (0.2 * (43 - 39.02)) = 39.82

LEAVE A LIKE FOR THE ANSWER IF YOU FOUND IT HELPFUL